Shift report 20011028 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Oct 28 2001 - 08:00:43 EST

  • Next message: Apache: "Shift report 20011028 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Oct 28 08:00:42 2001, Shift: 00:00-08:00
    Supervisor: JC, with: 
    Log book#10 pages 64 - 65
    
    Shift Summary:
    ------------------------------------------------------------
    1. 
    Two good stores. Data-taking almost all night. Shifting
    clock one hour back helped a little.
    
    2.
    Three data runs:
       				    FS		MRS	
                 Trig 1   Trig 6
      Run 5542    175k      55k	3deg, 1/3A   40deg, 3A	
      Run 5543     69k      22k     3deg, 1/3A   40deg, 3A
      Run 5544    386k     123k	3deg, 1/3A   40deg, 6A	
    
    Completed runs for:
     FS @  3deg, 1/3A - 201K trig-6 (runs 5542-5544)
    MRS @ 40deg,   3A - 422K trig-6 (runs 5525-5543)
    
    Present configuration:
     FS @  3deg, 1/3B -    0 trig-6 so far
    MRS @ 40deg,   6A - 123K trig-6 in run 5544
    
    Pedestals and DVM runs 5545-5547.
    
    3.
    No detector problems. I think. 
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Oct 28 2001 - 08:01:30 EST